### v1

# Pairwise Judgment Formulation for Semantic Embedding Model in Web Search

Abstract
Semantic Embedding Model (SEM), a neural network-based Siamese architecture, is gaining momentum in information retrieval and natural language processing. In order to train SEM in a supervised fashion for Web search, the search engine query log is typically utilized to automatically formulate pairwise judgments as training data. Despite the growing application of semantic embeddings in the search engine industry, little work has been done on formulating effective pairwise judgments for training SEM. In this paper, we make the first in-depth investigation of a wide range of strategies for generating pairwise judgments for SEM. An interesting (perhaps surprising) discovery reveals that the conventional pairwise judgment formulation strategy wildly used in the field of pairwise Learning-to-Rank (LTR) is not necessarily effective for training SEM. Through a large-scale empirical study based on query logs and click-through activities from a major commercial search engine, we demonstrate the effective strategies for SEM and highlight the advantages of a hybrid heuristic (i.e., Clicked > Non-Clicked) in comparison to the atomic heuristics (e.g., Clicked > Skipped) in LTR. We conclude with best practices for training SEM and offer promising insights for future research.
Keywords Search Engine  $\cdot$ Semantic Embedding Model  $\cdot$ Pairwise Judgment Formulation
1 Introduction
With the increasing awareness of the importance of latent semantics within text [14, 21, 23], the Semantic Embedding Model (SEM) attracts lots of attention from the information retrieval and natural language processing [4, 5, 28]. For the purpose of Web information retrieval, SEM is trained based on the search engine query log, which consists of queries, search results, and the record of a variety of user activities [15]. In practice, the queries and search results (i.e., the titles of the retrieved Web pages) are utilized to compose pairwise training instances. Formally, for a query $q$, if a title $p_{q}$ is preferred to a title $n_{q}$ by the user, they are formulated as a pairwise judgment $p_{q}$ > $n_{q}$, based on which SEM is trained in order to increase the similarity of ($q$, $p_{q}$) and the dissimilarity of ($q$, $n_{q}$).
A crucial issue of training SEM is to automatically formulate high-quality pairwise judgments, i.e., how to infer the user’s implicit preference based on the query log. The same problem is intensely studied in the field of pairwise Learning-to-Rank (LTR) [7, 20, 24]. Researchers propose an effective strategy for deriving pairwise judgment based on heuristics, which is inspired by the examination and click-through activities of users [11]. The core of these heuristics is to eliminate noise and position bias from the query log. Specifically, the heuristics proposed in [7, 20, 24] assume that a clicked document with the current query is preferred over an examined but not clicked document. With the intensive study of this topic in LTR, there exists little work on pairwise judgment formulation for embedding-based models such as SEM. Hence, it is interesting to study whether the well-established heuristics in LTR can be successfully applied to generate pairwise judgements for SEM.
In this paper, we investigate this problem based on the query log of a major commercial search engine and propose a series of strategies to formulate pairwise judgments. Through extensive experiments, we quantitatively compare these strategies and demonstrate some effective strategies for training SEM. An important insight obtained from such large-scale study is that the conventional pairwise judgment formulation heuristics in LTR are not necessarily effective for training SEM, which requires specialized strategies for deriving effective training data. We find that the strategy Clicked>Non-Examined which is rarely used in LTR works the best for SEM. The hybrid strategy Clicked>Non-Clicked is also demonstrated to have practical advantages in SEM. Such result reveals that investigating the strategies of pairwise judgment formulation is a valuable direction for practitioners who want to improve the performance of embedding-based models. The contributions of this paper are summarized as follows:
  1. 1.
    We make a detailed description of how to apply the SEM model in Web search scenarios;
  2. 2.
    We make the first in-depth study of formulating pairwise judgments for embedding-based models;
  3. 3.
    We make extensive experiments based on the largest (to the best of our knowledge) industrial-level pairwise judgement dataset and offer practical suggestions for training SEM.
The remainder of this paper is organized as follows. In Section 2, we review the most related work. In Section 3, we describe the background knowledge and experiment design. In Sections 4 and 5, we discuss the strategies for formulating pairwise judgments. Finally, we conclude this paper in Section 6.
2 Related Work
Formulating pairwise judgments is well-studied in pairwise LTR. Radlinksi et al. [24] proposed an approach to learning ranked retrieval functions by deriving pairwise judgment from a series of queries. Joachims et al. [20] examined the reliability of implicit feedback from click-through data in Web search. They show that preferences derived from user clicks are reasonably accurate, particularly when comparing results from the same query. The heuristics proposed in [20, 24] are widely used in conventional LTR scenarios.
Search engine users’ browsing patterns are also intensively studied in the field of click models. Chapelle et al. [2] consider click logs as an important source of implicit feedback. They proposed a Dynamic Bayesian Network which aims at providing unbiased estimation of the relevance from the click logs. A personalized click model is proposed in [25] to describe the user-oriented click preferences, which applies and extends tensor factorization from the view of collaborative filtering. [3] proposed a Noise-aware Clicked Model by characterizing the noise degree of a click. With their individual differences, almost all click models assume that search engine users sequentially browse the results from top to bottom and click some results if they are perceived as relevant. This assumption is strictly aligned with those studied in [20].
In recent years, learning semantic embedding has attracted lots of research attention. For example, Huang et al. [6] developed deep structured semantic model that projects queries and documents into a common low-dimensional space where the relevance of a document given a query is readily computed as the distance between them. The proposed model is discriminatively trained by maximizing the conditional likelihood of the clicked documents given a query using the click-through data. Shen et al. [26] developed a variant of the deep structured semantic model by incorporating a convolutional-pooling structure over word sequences to learn low-dimensional, semantic vector representations for search queries and Web documents. Recently, SEM has been proposed in [28] and is considered to be a more efficient model for learning semantic embedding. The major difference between SEM and its counterparts is the use of hinge loss with pairwise training paradigm rather than softmax-based loss functions. Hence, training SEM is typically more efficient since it does not need to conduct backpropagation for every training instance. With the success of SEM and its related models in the search engine industry [22, 27], there is surprisingly little work on studying the best approach for deriving training data for them. To the best of our knowledge, this work presents the pioneering effort to study the strategies of formulating pairwise judgments for training SEM.
3 Prerequisites
In Section 3.1, we discuss the basic knowledge of semantic embedding model. In Section 3.2, we introduce the terminology that is utilized for describing the pairwise judgments. In Section 3.3, we describe the dataset that is used in our experiments.
Figure 1: Semantic Embedding Model Architecture
3.1 Semantic Embedding Model for Web Search
In this subsection, we first elaborate the architecture of SEM and then we detail how to optimize its parameters.
Architecture
Inspired by the work in [28], we describe how the Semantic Embedding Model (SEM) works for information retrieval scenario. The architecture of SEM is shown in Figure 1. A hinge loss $L$ is used for training:
$L=\frac{1}{m}\sum_{i=1}^{m}\left(\cos\langle f(q_{i}),f(p_{q_{i}})\rangle-\cos\langle f(q_{i}),f(n_{q_{i}})\rangle\right)$ (1)
where $m$ is the amount of training instances, $\cos$ indicates cosine similarity and the function $f(\cdot)$ indicates the mapping from a query or a Web page title to a semantic embedding.
As shown in Figure 1, the bottom layer are the word embeddings. Through adding the embedding of words in query in an element-wise fashion, we obtain an intermediate representation of the query. If we denote $x$ as the input word embedding, $j=1,\dots,N$ as the term index of query, $h$ as the intermediate representation of query, and $i$ as the element id of $h$, then we have
$h_{i}=\sum_{j=0}^{N}x_{i}$ (2)
$g_{i}=softsign(h_{i})=\frac{h_{i}}{1+|h_{i}|}$ (3)
Then the intermediate representation goes through fully-connected neural network layers and is mapped to a final embedding. We denote $O$ as the final embedding, $W$ as the fully-connected neural network layer’s weight matrix, and $b$ as the bias, then we have
$O=Wh+b$ (4)
The two titles in a pairwise judgment are processing in analogous approach. Based on the final embeddings of the query and title that are denoted as $O_{q}$ and $O_{d}$, we calculate their cosine similarity as follows:
$\cos({O}_{q},{O}_{d})={\frac{{O}_{q}^{T}{O}_{d}}{\|{O}_{q}^{T}\||{O}_{d}\|}}$ (5)
When a query $q^{\prime}$ and its corresponding Web page titles come, we can calculate the semantic relevance between $q^{\prime}$ and the title of a Web page title $t^{\prime}$ by $\cos\langle f(q^{\prime}),f(t^{\prime})\rangle$, which can be directly utilized for ranking or work as a feature for sophisticated ranking algorithms.
Optimization
The neural network parameters and the word embeddings are updated by the conventional backpropagation. The SEM model is trained using stochastic gradient descent. Let $\Lambda$ be the parameters and $\Delta=\cos\langle f(q_{i}),f(p_{q_{i}})\rangle$, they are updated as follows:
$\Lambda_{t}=\Lambda_{t-1}-\gamma_{t}\frac{\partial\Delta}{\partial\Lambda_{t-1}}$ (6)
where $\Lambda_{t}$ and $\Lambda_{t-1}$ are the model parameters at $t^{th}$ iteration and $(t-1)^{th}$ iteration respectively. And $\gamma_{t}$ is the learning rate at $t^{th}$ iteration. This process is conducted for all training instances and repeated for several iterations until convergence is achieved. We derive the model parameters gradient as follows:
$\frac{\partial\Delta}{\partial\Lambda_{t-1}}=\frac{\partial\cos(O_{q},O_{d+})}{\partial\Lambda_{t-1}}-\frac{\partial\cos(O_{q},O_{d-})}{\partial\Lambda_{t-1}}$ (7)
To simplify the notation of calculating the derivatives of $W$, we let $d$ denote $d^{+}$ and $d^{-}$, and we let $a$, $b$, $c$ be $O^{T}_{q}O_{d}$, $\frac{1}{\|Q_{q}\|}$, and $\frac{1}{\|Q_{d}\|}$, respectively. Then, we can compute $\frac{\partial\Delta}{\partial W_{q}}$ and $\frac{\partial\Delta}{\partial W_{d}}$ by using the following formulas:
${\frac{\partial\mathrm{cos}({O}_{q},{O}_{d})}{\partial W_{q}}}={\frac{\partial}{\partial W_{q}}}{\frac{{O}_{q}^{T}{O}_{d}}{\|{O}_{q}\|\|{O}_{d}\|}}=\delta_{{O}_{q}}^{(q,d)}h_{d}^{T}$ (8)
${\frac{\partial\mathrm{cos}({O}_{q},{O}_{d})}{\partial W_{d}}}={\frac{\partial}{\partial W_{d}}}{\frac{{O}_{q}^{T}{O}_{d}}{\|{O}_{q}\|\|{O}_{d}\|}}=\delta_{{O}_{d}}^{(q,d)}h_{d}^{T}$ (9)
where $\delta_{{O}_{q}}^{(q,d)}=bcO_{d}-acb^{3}O_{q}$ and $\delta_{{O}_{d}}^{(q,d)}=bcO_{q}-acb^{3}O_{d}$
Similarly, we can compute the gradient of intermediate representation $\partial\Delta/\partial h$, then get the gradient of element-wise adding result $\partial\Delta/\partial v$. With softsign function in our model, each $\delta$ in the element-wise adding result can be calculated by:
$\delta_{v_{q}}^{(q,d)}=\frac{1}{(1+|V_{q}|)^{2}}\circ W_{q}^{T}\delta_{{O}_{q}}^{(q,d)}$ (10)
$\delta_{v_{d}}^{(q,d)}=\frac{1}{(1+|V_{d}|)^{2}}\circ W_{d}^{T}\delta_{O_{d}}^{(q,d)}$ (11)
In the above two formulas, the operator $\circ$ is the element-wise multiplication. Finally, we backpropagate the gradient of element-wise adding result $\delta_{v_{q}}^{(q,d)}$ and $\delta_{v_{d}}^{(q,d)}$ to each word embedding of query and documents.
3.2 Terminology
The Web search results of a query can be categorized into three groups based on the user’s click signals.
  1. 1.
    Clicked: The title of the result that is clicked by the users.
  2. 2.
    Skipped: The title of the result that is ranked above a clicked one, i.e., the result is examined by the user but no clicked.
  3. 3.
    Non-Examined: The title of the result that is ranked below all clicked ones.
3.3 Experimental Data
In our experiments, we gauge the performance by two metrics. The first metric is whether the model can effectively predict future clicking by scoring clicked results higher than non-clicked results. To achieve this goal, we prepare a testing dataset named Test-1 by deriving 23,000,000 pairwise judgments from the holdout query log. For a specific query and its top ten results, the pairwise judgment in Test-1 contains a randomly chosen clicked result and a randomly chosen non-clicked result. The second metric is to test whether the model can generate results that are aligned with human judgment. The second testing dataset named Test-2 is manually prepared by human experts and contains 530,000 pairwise judgments.
4 Atomic Strategies
In LTR, it is well believed that the relative preferences of clicked documents over skipped ones are reasonably accurate [1, 19]. Therefore, based on a user’s ranking preference reflected in his/her clicks, we propose several strategies to derive pairwise judgments. These strategies are mutually exclusive and can be used as the basic building blocks for deriving more complicated pairwise judgments.
  1. 1.
    Clicked > Skipped: This strategy is widely utilized in LTR applications. It assumes that the clicked results should be preferred to the skipped results.
  2. 2.
    Clicked > Clicked: This strategy differentiates clicked results by their click-through rate (CTR) and assumes that a result with higher CTR is preferred to that with lower CTR.
  3. 3.
    Clicked > Non-Examined: This strategy assumes that the clicked results are preferred to those that are not examined by the users.
  4. 4.
    Skipped > Non-Examined: This strategy is rarely applied in LTR since no click information is utilized. We include this strategy in our experiment for completeness.
Figure 2: Performance Comparison of Atomic Strategies on Test-1
Empirically, SEM usually takes several iterations to converge and we find that 50 iterations are typically enough to obtain a stable model. The experimental result on Test-1 is shown in Figure 2. We observe that Clicked>Non-Examined achieves the highest precision, showing that the pairwise judgments formed by Clicked>Non-Examined are of the best quality for training SEM. Pairwise judgments formed by Clicked>Skipped and Skipped>Non-Examined are of lower quality, and those formed by Clicked>Clicked are of the lowest quality.
We find that the result is quite different from the experience accumulated from conventional LTR. In LTR, the best pairwise judgment is typically formulated between skipped results and clicked results, i.e., the strategy Clicked>Skipped. However, for training SEM, the best strategy is Clicked>Non-Examined and the performance discrepancy between Clicked>Skipped and Clicked>Non-Examined is large. We further conduct an evaluation on Test-2 and the result is shown in Figure 3. We can see that the precision of the strategies on Test-2 is lower than those of Test-1, showing that the testing data in Test-2 is harder than those in Test-1. However, the trends of the lines in Test-2 and their relative positions are aligned with those in Figure 2.
Figure 3: Performance Comparison of Atomic Strategies on Test-2
Based on the results in Figures 2 and 3, we can see that Clicked>Clicked is the only strategy that makes the precision go down as more iterations are conducted, indicating that the pairwise judgments derived by this strategies is nearly useless. Another important observation is that the fluctuation of the precision of the other strategies: Clicked>Non-Examined has the least fluctuation, that of Clicked>Skipped is larger and that of Skipped>Non-Examined is the largest. The scale of fluctuation reveals the extent of the training distorted by occasionally encountering bad training instances.
5 Hybrid Strategy
From Figures 2 and 3, a large performance discrepancy between any two atomic strategies is observed. Intuitively, combining any two of these atomic strategies will not bring better results than Clicked>Non-Examined since the low-quality training instances will ”contaminate” the result. However, through extensive empirical evaluation, we find that the intuition holds but with one exception, which results in the following hybrid strategy: Clicked>Non-Clicked, which is a combination of Clicked>Skipped and Clicked>Non-Examined.
Figure 4: Performance Comparison of Hybrid Strategies on Test-1 and Test-2
Strategy Percentage
Clicked $>$ Skipped 22.25%
Clicked $>$ Clicked 5.96%
Clicked $>$ Non-Examined 38.85%
Skipped $>$ Non-Examined 32.92%
Table 1: Pairwise Judgment Distribution of Atomic Strategies
Based on the experimental results shown in Figure 4, we can see that Clicked>Non-Clicked slightly outperforms the best atomic strategy Clicked>Non-Examined on Test-1. In order to reveal the underlying reason, we present the distribution of the four atomic strategies. Based on the statistics shown in Table 1, we can see that no single strategy covers a majority of the pairs. Clicked>Skipped and Clicked>Non-Examined are responsible for 22.25% and 38.85% of the potential training data. Hence, relying on a single strategy such as Clicked>Non-Examined rules out many “fairly good” training instances and misses the chance of updating the embedding of many words. In the long run, the model using Clicked>Non-Clicked strategy sees more training instances achieves slightly better performance than the atomic strategies. The result shows that seeing more training instances is sometimes useful for training. Another possible explanation is that the testing data in Test-1 are essentially derived by Clicked>Non-Clicked strategy, hence, using Clicked>Non-Clicked has a better chance of getting a model with good fit for the testing data.
In order to see whether the performance of Clicked>Non-Clicked is exaggerated by Test-1, we present the experimental result of Test-2 in Figure 4. We can see that the results of Test-2 is different from that of Test-1. In the beginning, Clicked>Non-Examined performs better than Clicked>Non-Clicked, indicating that high-quality pairwise judgments give a better start for training SEM. In the long run, the performance discrepancy of the two strategies diminishes. Clicked>Non-Examined and Clicked>Non-Clicked achieves similar performance when convergence is achieved.
Based on the observations in Section 4 and Section 5. The insights we obtained from such large-scale study are summarized as follows:
  1. 1.
    The conventional LTR strategy for pairwise judgment formulation is not suitable for SEM. LTR aims to learn the weights of features in a ranking function. In contrast, the goal of SEM is to learn good representations of the features themselves.
  2. 2.
    The strategy Clicked>Non-Examined, which is rarely used in LTR, derives the best training data for SEM. Solely relying on this strategy significantly reduces the amount of training data for SEM and its performance is almost as good as using the strategy Clicked>Non-Clicked, which derives much more data for the training process to digest.
  3. 3.
    When Clicked>Non-Examined is applied, adding pairwise judgments derived from Clicked>Skipped to form a hybrid heuristic causes no harm for training SEM, and may slightly improve its performance by extending the diversity of training instances.
6 Conclusions
In this paper, we study the problem of formulating pairwise judgment for Semantic Embedding Model (SEM). We conduct an in-depth study of the information in the query log of a major search engine. Based on large-scale experiments, we quantitatively compare several strategies for pairwise judgment formulation. Experimental results verify the effectiveness of proposed strategies. More importantly, we reveal the fact that the pairwise judgment formulation for the neural network-based semantic embedding model is quite different from the pairwise LTR. Future work includes introducing more signals into the procedure of pairwise judgment formulation and investigating formulation strategies for SEM variants. The formulated pairwise judgment can also inspire downstream applications including topic discovery [9, 13, 16, 18], intent mining [10, 17], and user personalization [8, 12, 29].